OpenStudy (youngnephew34):

so x+2=y 2x=y+3 and you plug in x+2 in the 2x so2(x+2)=y+3 2x+4=y+3 how would you solve this problem

OpenStudy (anonymous):

that doesnt seem right, 2+x would not simplify to 2x. 2x is 2 times x

OpenStudy (anonymous):

http://www.mathpapa.com/calc.html?q=2x%20 @%20x%3D3

OpenStudy (youngnephew34):

im getting confuse

OpenStudy (youngnephew34):

how can you solve it @J-bird

OpenStudy (anonymous):

well, are you trying to solve for x or y??

OpenStudy (youngnephew34):

yes

OpenStudy (anonymous):

o,o

OpenStudy (anonymous):

i said x OR y

OpenStudy (youngnephew34):

butyou would have to plug in something @J-bird

OpenStudy (anonymous):

Hello, This is what we call a system of equations. Note your equations are: \(x+2=y\) \(2x=y+3\) To start to solve the system, choose which variable to solve for. In our case that would be \(x\) since one of the equations is set to \(y\). \(x+2=y\) \(\Rightarrow 2x=\color{green}{(x+2)}+3\) Do you follow so far?

OpenStudy (youngnephew34):

yes @Algorithmic

OpenStudy (anonymous):

So can you continue from where I left of at?

OpenStudy (youngnephew34):

i can try @Algorithmic

OpenStudy (anonymous):

Okay try, from where I left off.

OpenStudy (youngnephew34):

so you would have to put the 2 on the other side of the number @Algorithmic

OpenStudy (anonymous):

Well first let us condense (combine likes terms) in our equation. \[2x = x + 5\] Did you see how I did that?

OpenStudy (youngnephew34):

you added 3+2 right

OpenStudy (anonymous):

Correct, now to isolate variable \(x\). \[2x = x + 5\] \[\Rightarrow 2x - x = 5\] \[\Rightarrow x = 5\] Do you understand?

OpenStudy (youngnephew34):

i think i understand a lil bit @Algorithmic

OpenStudy (anonymous):

Okay, so we know now that \(x = 5\). We can now use that in our first equation to get \(y\). \[x + 2 = y\] \[\Rightarrow 5 + 2 = y\]

OpenStudy (youngnephew34):

so plug in 5 in the x

OpenStudy (anonymous):

Correct, and you get \(y\). So now you have both the values for \(x\) and \(y\).

OpenStudy (youngnephew34):

thanks a lot and do you have a kik

OpenStudy (anonymous):

You are welcome. No I do not have a Kik.

OpenStudy (youngnephew34):

ok